The measure of the complementary angle and the supplementary angle of a given angle is x and 4x respectively. What is the mearue of the given angle?
Answer by Theo(13342)

let y be the angle.
you get:
90 - y = x
180 - y = 4x
since x = 90 - y, you can substitute in the second equation to et:
180 - y = 4(90 - y)
simplify to get:
180 - y = 360 - 4y
subtract 180 from both sides of this equation and add 4y to both sides of this equation to get:
4y - y = 360 - 180
simplify to get;
3y = 180
divide both sides of this equation to get:
y = 60
that's your answer.
90 - y = = 90 - 60 = 30
180 - y = 180 - 60 = 120
120 is equal to 4 * 30.
this confirms the answer is good.